OpenStudy (anonymous):

Solve the equation below for x using any method: 2x^2+3x+6=0 A. (-3±i√39)/4 B. 2±3i C. x=-3;x=2 D. x=3±√57

OpenStudy (anonymous):

Hello Kassia : )

OpenStudy (anonymous):

hello :)

OpenStudy (anonymous):

Hi Kassia! I am back!

OpenStudy (anonymous):

A. (-3±i√39)/4

OpenStudy (anonymous):

2x^2 + 3x + 6 = 0 for this purpose is best solved using the quadratic formula. -b +- sqrt(b^2 - 4ac)/2a. a = 2 b = 3, c = 6. The answer is A. Hope this helps. If it does, please choose good answer. Visit my website at http://www.tutorsean.net for free math training videos.

OpenStudy (anonymous):

\[x = \frac{ -3 \pm \sqrt{9 - 4*2*6}}{2}\] \[x = \frac{ -2 \pm i\sqrt{39}}{2}\] A

OpenStudy (anonymous):

sorry 3 not 2 typo

OpenStudy (anonymous):

and 4 in denominator 2*2 = 4 lot of typo sorry